# GARG-AML against Smurfing: A Scalable and Interpretable Graph-Based Framework for Anti-Money Laundering

### v2

@@ 4.3 Evaluation
Two key aspects are critical for business adoption. The first is model performance—a robust model must accurately detect money laundering while keeping false positives to a minimum. The second is scalability. Real-world transaction datasets are massive, with banks processing thousands to millions of transactions daily. Therefore, any proposed method must scale efficiently to such volumes. Accordingly, the evaluation of the methods focuses on these two dimensions.
For performance evaluation, the GARG-AML scores are assessed against ground-truth labels. The evaluation is conducted using the area under the receiver operating characteristic curve (AUC-ROC) and the area under the precision-recall curve (AUC-PR). These scores are also used to rank the methods, where the best-performing method receives rank 1, the second-best method receives rank 2, and so on.
While AUC-ROC is a widely used metric for binary classification, it may not be sufficient for anti-money laundering (AML) datasets, which are often highly imbalanced—typically with fewer than $0.1\%$ observations flagged as laundering money. To address this imbalance, we include AUC-PR, which has been shown to be a more reliable performance metric under situations of extreme label imbalance (Davis & Goadrich, 2006; Ozenne et al., 2015). In general terms, the AUC-PR is the expected average precision when a random recall threshold is selected. In other words, the AUC-PR measures the model’s ability to identify positive examples correctly while minimising false positives across all possible threshold values. The AUC-PR specifically focuses on the model’s precision in identifying the positive class at various recall thresholds.
Scalability is assessed in terms of calculation time. In addition to runtime, memory usage is also a key factor. We also report any instances where methods fail due to memory limitations. The theoretical time complexities are $O\left(k\mid E\mid\log\left(\mid V\mid\right)\right)$ for Flowscope (Li et al., 2020) and $O\left(\mid V\mid\cdot\mid E\mid+\mid V\mid^{2}\right)$ for AutoAudit (Lee et al., 2020).
For GARG-AML, the main calculations for each individual node are the selection of the second-order neighbourhood and the assignment of the neighbours to the levels (in the directed case). Assuming that we are working in sparse graphs, both steps have complexity $O\left(\mid V\mid\right)$. This brings us to a total worst case time complexity of $O\left(n^{2}\right)$. Hence, theoretically, GARG-AML scales as AutoAudit, while Flowscope is much faster. However, we can speed up the calculations by doing running them in parallel. In addition, second-order neighbours are generally much smaller than $n$, meaning that also those steps can run very quickly.
@@ 4.4 Statistical Test
The statistical significance of the difference in ranks is evaluated using the Friedman test (Friedman, 1937, 1940) and the post hoc Nemenyi test (Nemenyi, 1963), as described by Demšar (2006). It compares the performance of the $k$ different methods across the $N$ datasets. If we take $r_{i}^{j}$ to be the rank of method $j$ on dataset $i$ for some performance metric, then the average performance across all datasets is equal to $R_{j}=\frac{1}{N}\sum_{i}r_{i}^{j}$. In our experiments, the Friedman and Nemenyi statistical tests compare the ranks of the 8 $(k)$ methods across the 66 $(N)$ synthetic datasets. We apply a significance level of $0.05$ for these tests.
@@
Rejecting the null hypothesis indicates that significant differences are present among the methods. We then proceed with the post hoc Nemenyi test, which finds that the performances of two methods are significantly different at level $\alpha$ if their respective average ranks differ by at least the critical difference (CD):
@@
where critical levels $q_{\alpha}$ are provided by Demšar (2006). We visualize the results using CD diagrams (Demšar, 2006) for the AUC-ROC and AUC-PR.
@@ 5 Results and Discussion
The results are presented in two parts. First, we present and discuss the scalability of the methods. Afterward, we show the results for the performance and the CD diagrams.
@@ 5.1 Scalability
Figure 7 and Figure 7 hows the calculation times across the various network sizes of the synthetic data and IBM datasets, respectively. The time in Figure 7 is shown on a logarithmic scale since this provides a better distinction among the methods across dataset sizes. These results indicate that FlowScope is the fastest of all methods, although GARG-AML seems to also scale well to larger datasets, contrary to what we expected from the theoretical worst-case time complexity. In contrast, AutoAudit—although quite fast in the beginning—has difficulties with networks of the largest size. In addition, the distribution of the calculation time of AutoAudit is much wider and has more outliers, which indicates that it is highly sensitive to the network topology.
Figure 6: Box plot of the run time for each method on a logarithmic scale over the three different sizes of synthetic datasets. From left to right are FlowScope, AutoAudit, undirected GARG-AML and directed GARG-AML scores.
Figure 7: Bar plot of the run time for each method on a logarithmic scale over the two IBM datasets. From left to right are FlowScope, undirected GARG-AML and directed GARG-AML scores.
Another key observation is that AutoAudit consistently runs out of time—exceeding 16 hours—on Watts-Strogatz networks of size 10,000. This is reflected by setting the out-of-time results equal to 16 hours (57,600 seconds) in Figure 7, which results in outliers for AutoAudit. For the larger networks, i.e. the synthetic network of size 100,000 and the IBM datasets, AutoAudit exceeds the available memory. Hence, the results for AutoAudit are omitted entirely from Figure 7.
@@ 5.2.1 Synthetic Data
The results are presented according to the different injected patterns. These are provided in Figure 8, with their corresponding ranks presented in Figure 9.
Figure 8: AUC-ROC and AUC-PR values of the compared methods across all synthetic datasets, with the split-on money laundering pattern.
Figure 9: Ranks of the AUC-ROC and AUC-PR values of the compared methods across all synthetic datasets, with the split-on money laundering pattern, where rank 1 indicates the best performance and rank 8 indicates the worst performance.
The AUC-PR values are more spread out than the AUC-ROC values. This is caused by the varying level of class imbalance, as shown in Table 4. For a model with random label predictions, the AUC-ROC is always 0.5, regardless of the label distribution, whereas the AUC-PR equals the proportion of positive labels, which leads to the high variance shown in Figure 8.
Figure 8 indicates that, on average, the undirected GARG-AML scores consistently showing great performance. This performance is boosted when combining the GARG-AML scores with the tree-based learners. In comparison, the performance of AutoAudit is less stable. Here, we set all the AUC-ROC and AUC-PR values to 0 when the method runs out of time or out of memory to allow for a full comparison of the ranks. For these cases, AutoAudit ranks last compared with the other methods.
The ranks in Figure 9 corroborate the results from before. We can see more clearly that combining the tree-based learners with the scores and summary statistics of the node’s neighbours score better than the other methods, even when just applying a simple decision tree. We test the statistical significance of the difference in average rank using the Friedman test. For all the cases, the Friedman test rejects the null hypothesis, which indicates that significant differences between the ranks are present.
@@
The post hoc Nemenyi test is used to discern which methods have significantly different average ranks. The results are shown in Figure 11 and Figure 11 for the AUC-ROC and AUC-PR, respectively. Here, again the tree-based methods differ significantly from most other methods, and the undirected scores on their own also obtain strong performance. For the existing mules patterns, however, it seems that the best performing models do not differ significantly from FlowScope.
@@
Figure 10: Critical distance plots of the average ranks according to the AUC-ROC for the considered patterns.
@@
Figure 11: Critical distance plots of the average ranks according to the AUC-PR for the considered patterns.
@@ 5.2.2 IBM Data
This section presents the results on the IBM datasets for the smurfing patterns, i.e., gather-scatter and scatter-gather, and the full ‘Is Laundering’ label. The thresholds applied for binary label definition are 0.1, 0.2 and 0.9. The results for the HI-Small and LI-Large datasets are presented in Figure 12 and Figure 13, respectively. The results for all patterns and thresholds are provided in Appendix B.
Figure 12: The AUC-ROC (solid bars) and AUC-PR (hatched bars) results on the HI-Small IBM dataset on the money laundering and smurfing labels across different thresholds.
Figure 13: The AUC-ROC (solid bars) and AUC-PR (hatched bars) results on the LI-Large IBM dataset on the money laundering and smurfing labels across different thresholds. Some panels are empty, due to the absence of positive labels for these cases.
@@
Similar to the results on the synthetic data, the undirected version of the GARG-AML scores obtain best performance for almost all cases for the AUC-ROC values, followed by the tree-based models. For both datasets, the undirected GARG-AML scores give great results for the general laundering cases as well, which include more patterns than just smurfing.
@@
For the lower threshold values, we see a significant improvement of AUC-PR across the tree-based methods. This results from these methods being able to better capture the interaction among a node’s GARG-AML score and those of its neighbours. This reduces the false positives compared to the basic score, resulting in a much higher AUC-PR. Therefore, on lower thresholds, we conclude that the tree-based methods are better suited for smurfing detection.
@@
For the highest thresholds the basic GARG-AML scores provide great performance, with AUC-ROC above $80\%$. In these cases, only clients are labelled that have $90\%$ or more of their transactions involved in money laundering. Therefore, low levels of obfuscation are present and the transaction patterns are very close to pure smurfing, making them easily recognisable by GARG-AML.
@@
The tree-based models, on the other hand, seem to perform poorly, in line with FlowScope. It seems that the class imbalance (see also Table 7 and Table 8) is too extreme in these cases, resulting in the model’s inability to train properly.

### v3

@@ 4.3 Evaluation
Two aspects are critical for business adoption. A model must (1) catch launderers with few false positives (model performance), and (2) process millions of daily transactions quickly (scalability). Accordingly, the evaluation of the methods focuses on these two dimensions.
The performance is evaluated using precision, F1-score, the area under the receiver operating characteristic curve (AUC-ROC) and the area under the precision-recall curve (AUC-PR).
[…]
Both AUC metrics, on the other hand, compare models across classification thresholds. The AUC-ROC quantifies the model’s ability to discriminate between classes by plotting the true positive rate against the false positive rate. While it is a widely used metric for binary classification, it may not be sufficient for anti-money laundering (AML) datasets, which are often highly imbalanced—typically with fewer than $0.1\%$ observations flagged as laundering money.
The AUC-PR is more indicative of performance under extreme class imbalance, where the positive class (smurfing) is rare (Davis and Goadrich, 2006; Ozenne et al., 2015). It plots precision against recall. The AUC-PR measures the model’s ability to identify positive examples correctly while minimising false positives across all possible threshold values. The AUC-PR specifically focuses on the model’s precision in identifying the positive class at various recall thresholds.
Scalability is assessed in terms of calculation time. The theoretical time complexity for the baseline models is $O\left(k\mid E\mid\log\left(\mid V\mid\right)\right)$ for Flowscope (Li et al., 2020) and $O\left(\mid V\mid\cdot\mid E\mid+\mid V\mid^{2}\right)$ for AutoAudit (Lee et al., 2020).
For GARG-AML, the main calculations are the selection of the second-order neighbourhood and the assignment of the neighbours to the levels. Assuming that we are working in sparse graphs, both steps have complexity $O\left(\mid V\mid\right)$. This brings us to a total worst case time complexity of $O\left(n^{2}\right)$. Hence, theoretically, GARG-AML scales as AutoAudit, while Flowscope is much faster. However, we can speed up the calculations by running them in parallel. In addition, second-order neighbours are generally much smaller than $n$, meaning that those steps can run very quickly.
In addition to runtime, memory usage is a key factor. We report any instances where methods fail due to memory limitations.
@@ 4.4 Statistical Test
To test whether the observed differences in performance are not due to random chance, we determine for each model its rank compared to the other models based on the AUC-ROC and the AUC-PR. The best-performing method receives rank 1, the second-best method receives rank 2, and so on.
The statistical significance of the difference in ranks is evaluated using the Friedman test (Friedman, 1937, 1940) and the post hoc Nemenyi test (Nemenyi, 1963), as described by Demšar (2006). It compares the performance of $k$ methods across $N$ datasets. If we take $r_{i}^{j}$ to be the rank of method $j$ on dataset $i$, the average performance across all datasets is equal to $R_{j}=\frac{1}{N}\sum_{i}r_{i}^{j}$.
@@
Rejecting the null hypothesis indicates that significant differences are present among the methods. In that case, we proceed with the post hoc Nemenyi test. It indicates that the performances of two methods are significantly different at level $\alpha$ if their respective average ranks differ by at least the critical difference (CD):
@@
where critical levels $q_{\alpha}$ are provided by Demšar (2006).
Our experiments compare the ranks of the $k=8$ methods across the $N=66$ synthetic datasets at significance level of $\alpha=0.05$. We visualize the results using CD diagrams (Demšar, 2006) for the AUC-ROC and AUC-PR.
@@ 5.1 Scalability
Figure 7 and Figure 7 show the calculation times across the various network sizes of the synthetic and IBM datasets, respectively. The time in Figure 7 is shown on a logarithmic scale, to provide a better distinction across dataset sizes. These results indicate that FlowScope is the fastest method, although GARG-AML seems to scale well to larger datasets, contrary to what we expected from the theoretical worst-case time complexity. In contrast, AutoAudit—although quite fast in the beginning—has difficulties with the largest networks. In addition, the distribution of the calculation time of AutoAudit is much wider with more outliers, indicating a high sensitivity to the network topology.
Figure 6: Box plot of the run time on a logarithmic scale across network sizes. From left to right: FlowScope, AutoAudit, undirected GARG-AML and directed GARG-AML scores.
Figure 7: Bar plot of the run time on a logarithmic scale for the two IBM datasets. From left to right: FlowScope, undirected GARG-AML and directed GARG-AML scores.
Additionally, AutoAudit consistently runs out of time—exceeding 16 hours—on Watts-Strogatz networks of size 10,000. We set the out-of-time results equal to 16 hours (57,600 seconds) in Figure 7, resulting in the observed outliers. For the larger networks, i.e. the synthetic network of size 100,000 and the IBM datasets, AutoAudit exceeds the available memory. Hence, AutoAudit is omitted from Figure 7.
@@ 5.2.1 Synthetic Data
The results are presented according to the injected patterns in Table 9, with more detailed results given in Appendix B. The results indicate that, on average, the undirected GARG-AML scores consistently show great performance. This performance is boosted when combining the GARG-AML scores with the tree-based learners. In comparison, the performance of the baselines, Flowscope and AutoAudit, is less stable.
A major concern for transaction monitoring is the high false positive rate. The precision shows that more than half of the alerts from FlowScope and AutoAudit were false positives. GARG-AML scores combined with tree-based learners result in more than $70\%$ of alerts being actual laundering accounts, except for the ‘existing mules’ pattern. This points to a strong efficiency gain for financial institutions.
Table 9: Performance Comparison Across Patterns
pattern method Precision F1 AUC-ROC AUC-PR
laundering flowscope $0.417\pm 0.348$ $0.192\pm 0.182$ $0.547\pm 0.097$ $0.272\pm 0.290$
autoaudit $0.219\pm 0.312$ $0.194\pm 0.268$ $0.374\pm 0.314$ $0.336\pm 0.334$
gargaml undirected $0.539\pm 0.458$ $0.460\pm 0.365$ $0.723\pm 0.128$ $0.426\pm 0.371$
gargaml directed $0.340\pm 0.365$ $0.164\pm 0.176$ $0.491\pm 0.273$ $0.265\pm 0.342$
gargaml tree undirected $\boldsymbol{0.929\pm 0.087}$ $0.801\pm 0.090$ $0.837\pm 0.080$ $0.707\pm 0.143$
gargaml boost undirected $0.886\pm 0.170$ $\boldsymbol{0.803\pm 0.148}$ $\boldsymbol{0.854\pm 0.099}$ $\boldsymbol{0.718\pm 0.193}$
gargaml tree directed $0.861\pm 0.183$ $0.720\pm 0.185$ $0.791\pm 0.110$ $0.615\pm 0.224$
gargaml boost directed $0.841\pm 0.220$ $0.728\pm 0.220$ $0.811\pm 0.112$ $0.631\pm 0.265$
separate flowscope $0.174\pm 0.313$ $0.165\pm 0.286$ $0.584\pm 0.200$ $0.196\pm 0.274$
autoaudit $0.113\pm 0.171$ $0.158\pm 0.236$ $0.509\pm 0.442$ $0.453\pm 0.447$
gargaml undirected $0.278\pm 0.261$ $0.370\pm 0.322$ $0.872\pm 0.115$ $0.363\pm 0.369$
gargaml directed $108\pm 0.166 0.0$$077\pm 0.081$ $0.754\pm 0.108$ $0.159\pm 0.143$
gargaml tree undirected $0.860\pm 0.240$ $0.841\pm 0.240$ $0.907\pm 0.140$ $0.795\pm 0.271$
gargaml boost undirected $\boldsymbol{0.896\pm 0.201}$ $\boldsymbol{0.875\pm 0.196}$ $\boldsymbol{0.930\pm 0.114}$ $\boldsymbol{0.828\pm 0.240}$
gargaml tree directed $0.845\pm 0.272$ $0.687\pm 0.247$ $0.802\pm 0.130$ $0.575\pm 0.259$
gargaml boost directed $0.762\pm 0.279$ $0.726\pm 0.270$ $0.873\pm 0.140$ $0.620\pm 0.310$
existing mules flowscope $0.223\pm 0.203$ $0.206\pm 0.196$ $0.596\pm 0.114$ $0.152\pm 0.152$
autoaudit $062\pm 0.094 0.0$$087\pm 0.129$ $318\pm 0.271 0.0$$071\pm 0.090$
gargaml undirected $034\pm 0.079 0.0$$053\pm 0.119$ $396\pm 0.159 0.0$$059\pm 0.089$
gargaml directed $0.120\pm 0.166$ $0.138\pm 0.188$ $0.588\pm 0.117$ $0.108\pm 0.159$
gargaml tree undirected $0.337\pm 0.367$ $0.281\pm 0.312$ $0.621\pm 0.150$ $0.225\pm 0.247$
gargaml boost undirected $\boldsymbol{0.373\pm 0.360}$ $\boldsymbol{0.325\pm 0.304}$ $\boldsymbol{0.643\pm 0.154}$ $\boldsymbol{0.244\pm 0.238}$
gargaml tree directed $0.260\pm 0.338$ $0.194\pm 0.249$ $0.574\pm 0.106$ $0.152\pm 0.178$
gargaml boost directed $0.323\pm 0.311$ $0.263\pm 0.257$ $0.614\pm 0.135$ $0.185\pm 0.185$
new mules flowscope $020\pm 0.051 0.0$$014\pm 0.030$ $458\pm 0.066 0.0$$069\pm 0.098$
autoaudit $043\pm 0.072 0.0$$059\pm 0.094$ $241\pm 0.221 0.0$$056\pm 0.082$
gargaml undirected $0.227\pm 0.207$ $0.306\pm 0.268$ $0.756\pm 0.181$ $0.197\pm 0.185$
gargaml directed $112\pm 0.137 0.0$$096\pm 0.109$ $401\pm 0.136 0.0$$088\pm 0.116$
gargaml tree undirected $0.706\pm 0.320$ $0.580\pm 0.274$ $0.749\pm 0.141$ $0.466\pm 0.242$
gargaml boost undirected $0.702\pm 0.276$ $0.657\pm 0.260$ $0.835\pm 0.152$ $0.551\pm 0.263$
gargaml tree directed $0.680\pm 0.332$ $0.607\pm 0.281$ $0.778\pm 0.141$ $0.482\pm 0.264$
gargaml boost directed $\boldsymbol{0.732\pm 0.276}$ $\boldsymbol{0.719\pm 0.265}$ $\boldsymbol{0.863\pm 0.159}$ $\boldsymbol{0.607\pm 0.305}$
These results are an average over widely different label distributions (see Table 4). The precision, F1-scores and the AUC-PR are highly sensitive to the label imbalance. Their baselines depend on the ratio of positive labels to the total number of labels. This changes with each dataset. The AUC-ROC baseline on the other hand is always equal to 0.5. This makes all but the AUC-ROC values in Table 9 hard to compare.
To mitigate this, we analyse the average rank of each model’s AUC-ROC and AUC-PR for a more robust evaluation. The results are shown in Figure 9 and Figure 9 for the AUC-ROC and AUC-PR, respectively. We find that the tree-based learners score better than the other methods, even when applying a simple decision tree.
@@
For all cases, the Friedman test rejects the null hypothesis, which indicates that significant differences between the ranks are present. The post hoc Nemenyi indicates that the tree-based methods differ significantly from most other methods. The undirected scores on their own already obtain strong performance. For the existing mules patterns, however, it seems that the best performing models do not differ significantly from FlowScope.
@@
Figure 8: Critical distance plots of the average ranks according to the AUC-ROC across patterns.
@@
Figure 9: Critical distance plots of the average ranks according to the AUC-PR across patterns.
@@ 5.2.2 IBM Data
Table 10 and Table 11 present the results for the smurfing patterns, i.e., gather-scatter and scatter-gather, and the full ‘Is Laundering’ label. The cut-offs for binary label definition are 0.1, 0.5 and 0.9. The results for all patterns and cut-offs are provided in Appendix C. Missing results indicate that there were too few labels for the model to learn from. The tree-based models could not be trained on the HI-Small dataset with cut-off 0.9 (Table 10). No positive Gather-Scatter and Scatter-Gather labels were present in the LI-Large dataset for cut-off 0.9 (Table 11).
Table 10: Performance comparison for the HI-Small dataset.
Pattern Model Cutoff 0.1 Cutoff 0.5 Cutoff 0.9
Precision F1 AUC-ROC AUC-PR Precision F1 AUC-ROC AUC-PR Precision F1 AUC-ROC AUC-PR
Is Laundering FlowScope 0.000 0.000 50.0 0.445 0.000 0.000 50.0 0.020 0.000 0.000 50.0 0.013
GARG-AML Undirected 0.006 0.012 61.2 0.559 0.000 0.001 66.2 0.037 0.000 0.001 81.7 0.041
GARG-AML Directed 0.003 0.005 41.2 0.342 0.000 0.000 46.0 0.018 0.000 0.000 53.3 0.015
GARG-AML Undir. Tree 0.476 0.055 51.4 1.549 0.000 0.000 50.0 0.020 0.000 0.000 50.0 0.014
GARG-AML Undir. Boost 0.316 0.017 50.7 0.958 0.077 0.045 50.0 0.020 0.000 0.000 50.0 0.014
GARG-AML Dir. Tree 0.543 0.053 51.4 1.932 0.000 0.000 50.0 0.020 0.000 0.000 50.0 0.014
GARG-AML Dir. Boost 0.576 0.053 51.4 2.023 0.000 0.000 50.0 0.020 0.000 0.000 50.0 0.014
Gather-Scatter FlowScope 0.000 0.000 50.0 0.056 0.000 0.000 50.0 0.003 0.000 0.000 50.0 0.001
GARG-AML Undirected 0.001 0.002 65.7 0.082 0.000 0.000 53.5 0.003 0.000 0.000 83.4 0.004
GARG-AML Directed 0.000 0.001 46.6 0.049 0.000 0.000 47.8 0.003 0.000 0.000 63.9 0.002
GARG-AML Undir. Tree 0.200 0.022 50.6 0.346 0.000 0.000 50.0 0.003 0.000 0.000 50.0 0.001
GARG-AML Undir. Boost 0.107 0.053 52.3 0.391 0.000 0.000 50.0 0.003 0.000 0.000 50.0 0.001
GARG-AML Dir. Tree 0.000 0.000 50.0 0.056 0.000 0.000 50.0 0.003 0.000 0.000 50.0 0.001
GARG-AML Dir. Boost 0.098 0.063 52.3 0.507 0.000 0.000 50.0 0.003 0.000 0.000 50.0 0.001
Scatter-Gather FlowScope 0.000 0.000 50.0 0.036 0.000 0.000 50.0 0.003 0.000 0.000 50.0 0.000
GARG-AML Undirected 0.001 0.001 57.3 0.040 0.000 0.000 41.3 0.002 0.000 0.000 86.0 0.001
GARG-AML Directed 0.000 0.000 41.4 0.029 0.000 0.000 31.2 0.002 0.000 0.000 46.9 0.000
GARG-AML Undir. Tree 0.486 0.387 66.1 17.557 0.000 0.000 50.0 0.003 - - - -
GARG-AML Undir. Boost 0.274 0.288 66.1 10.544 0.043 0.080 62.5 0.660 - - - -
GARG-AML Dir. Tree 0.529 0.247 58.0 8.539 0.000 0.000 50.0 0.003 - - - -
GARG-AML Dir. Boost 0.364 0.270 60.7 7.821 0.034 0.061 62.5 0.864 - - - -
Table 11: Performance comparison for the LI-Large dataset.
Pattern Model Cutoff 0.1 Cutoff 0.5 Cutoff 0.9
Precision F1 AUC-ROC AUC-PR Precision F1 AUC-ROC AUC-PR Precision F1 AUC-ROC AUC-PR
Is Laundering FlowScope 0.000 0.000 50.0 0.216 0.000 0.000 50.0 0.002 0.000 0.000 50.0 0.000
GARG-AML Undirected 0.003 0.006 55.1 0.255 0.000 0.000 68.0 0.003 0.000 0.000 91.3 0.003
GARG-AML Directed 0.002 0.004 43.9 0.178 0.000 0.000 62.4 0.002 0.000 0.000 72.1 0.001
GARG-AML Undir. Tree 0.000 0.000 50.0 0.216 0.000 0.000 50.0 0.002 0.000 0.000 50.0 0.000
GARG-AML Undir. Boost 0.000 0.000 50.1 0.226 0.000 0.000 50.0 0.002 0.000 0.000 50.0 0.000
GARG-AML Dir. Tree 0.333 0.001 50.0 0.241 0.000 0.000 50.0 0.002 0.000 0.000 50.0 0.000
GARG-AML Dir. Boost 0.000 0.000 50.0 0.216 0.000 0.000 53.8 0.002 0.000 0.000 50.0 0.000
Gather-Scatter FlowScope 0.000 0.000 50.0 0.008 0.000 0.000 50.0 0.000 - - - -
GARG-AML Undirected 0.000 0.000 59.5 0.009 0.000 0.000 59.0 0.001 - - - -
GARG-AML Directed 0.000 0.000 56.6 0.008 0.000 0.000 58.8 0.000 - - - -
GARG-AML Undir. Tree 0.000 0.000 50.0 0.008 0.000 0.000 50.0 0.000 - - - -
GARG-AML Undir. Boost 0.000 0.000 49.9 0.008 0.000 0.000 63.1 0.001 - - - -
GARG-AML Dir. Tree 0.000 0.000 50.0 0.007 0.000 0.000 50.0 0.000 - - - -
GARG-AML Dir. Boost 0.000 0.000 50.0 0.007 0.000 0.000 50.0 0.000 - - - -
Scatter-Gather FlowScope 0.000 0.000 50.0 0.015 0.000 0.000 50.0 0.000 - - - -
GARG-AML Undirected 0.000 0.000 58.8 0.018 0.000 0.000 41.6 0.000 - - - -
GARG-AML Directed 0.000 0.000 56.1 0.015 0.000 0.000 40.7 0.000 - - - -
GARG-AML Undir. Tree 0.000 0.000 50.0 0.015 0.000 0.000 50.0 0.000 - - - -
GARG-AML Undir. Boost 0.000 0.000 50.0 0.015 0.000 0.000 49.9 0.000 - - - -
GARG-AML Dir. Tree 0.000 0.000 50.0 0.015 0.000 0.000 50.0 0.000 - - - -
GARG-AML Dir. Boost 0.000 0.000 50.0 0.015 0.000 0.000 50.0 0.000 - - - -
The precision and F1-scores are all (close) to zero. This is a result of the choice of the widely-used threshold value of 0.5. Almost no observations get a score of 0.5 or higher, meaning that almost all observations are classified as normal. This highlights that careful tuning is needed for these thresholds, but also that comparing models on just a few thresholds can be problematic. Therefore, we decide to rely mostly on the threshold-independent metrics (AUC-ROC and AUC-PR) to form our conclusions.
@@
Similar to the results on the synthetic data of Section 5.2.1, the undirected version of the GARG-AML scores obtains best performance for almost all cases for the AUC-ROC values, followed by the tree-based models. The undirected GARG-AML scores give great results for the general laundering cases, which include more patterns than just smurfing.
@@
For the lower cut-off values, we see a significant improvement of AUC-PR across the tree-based models. This results from the models being able to better capture the interaction among a node’s GARG-AML score and those of its neighbours. This reduces the false positives compared to the base score, resulting in a much higher AUC-PR. Therefore, on lower cut-offs, we conclude that the tree-based methods are better suited for smurfing detection.
@@
For the highest cut-offs, the base GARG-AML scores provide great performance, with AUC-ROC above $80\%$. In these cases, only clients are labelled that have $90\%$ or more of their transactions involved in money laundering. Therefore, these patterns closely resemble pure smurfing, making them easily recognisable by GARG-AML.
@@
The tree-based models, on the other hand, seem to perform poorly, in line with FlowScope. The class imbalance (see also Table 7 and Table 8) is too extreme in these cases, resulting in the model’s inability to train properly.